Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: Last week of CnB (B configuration move begins on Monday); severe weather over the next day+; RSRO still seeing missing sub-bands. Proposals due today at 1500 MST!
    • News from the ALMA Computing Planning Meeting (Joe/Jeff)
  • Key dates: * 01 Feb: Proposal deadline for D configuration+ * 03-04 Feb: Dish verification review ($KA)
  • Staff Travel:
  • Visitors:
    • 05 Jan - 06 Apr: John Cannon sabbatical at DSOC

Commissioning priorities for the week (McMullin, Chandler)

Single maintenance day week:
  • Mon-Tue, Thu-Fri:
    • 0900-1700 Testing/software development time
    • 1800-0900: Science operations
  • Wed: Maintenance day

Top Issues:
  • Correlator: fswitch sub-band anti-aliasing causing phantom spectral lines when using recirculation
  • Correlator: missing sub-bands (initially thought tied to the station board upgrades; backed out); still occurring through 29 Jan.
  • Correlator: time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags.
  • Construction: RFI sweep at Ku, K, Ka, Q pre-requisite for high pass filters for use with L301.
  • Array: Switched power heuristics for reduction.
  • Array/Construction: Dhawan/Rupen: RF switch tests with rapid monitoring.

Hardware status and updates (Langley, Perley, Rupen)

Software status and updates (Butler)

Commissioning Reports

Spectral Line: Goss/Chomiuk: Bandpass stability monitoring

  • Main event: report on bandpass stability monitoring; results posted at ECSV-107

Commissioning Group updates and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!

RSRO/ECSO Reports

Individual reports (updates)

Misc:

Demo Science Status

AOB

  • Action: Miriam: Review priorities with Jeff to see if integration of autoflag can be raised for this cycle. Discussion begun.
  • Action: Joe: Post Urvashi's autoflag test information - get feedback for next meeting.Details on using the LOFAR package for autoflag are posted at: https://bugs.nrao.edu/browse/CAS-2676

Misc Actions

  • [Frail/Butler]: Enable restrictions based on proximity to sun; review OST weather trending.
  • [McMullin/Frail]: %REDWrite up disposition of RR items.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • Development of WIDAR guidelines document (Michael)
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.

-- JosephMcMullin - 2011-01-31
Topic revision: r2 - 2011-02-01, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback